OpenStudy (anonymous):

what is the sum of 3/2x and 4/3x expressed in simplest form?

jimthompson5910 (jim_thompson5910):

is 3/2x really \[\Large \frac{3}{2}x\] or is it \[\Large \frac{3}{2x}\] ???

OpenStudy (anonymous):

it is 3 over 2x the 2nd choice

jimthompson5910 (jim_thompson5910):

thanks

jimthompson5910 (jim_thompson5910):

In order to add these fractions, we need the denominators to be the same

jimthompson5910 (jim_thompson5910):

So how do we do that?

jimthompson5910 (jim_thompson5910):

Any ideas?

OpenStudy (anonymous):

is the answer 7/5x

jimthompson5910 (jim_thompson5910):

not quite, you cannot just add the denominators

OpenStudy (anonymous):

o then i dont know what to do

jimthompson5910 (jim_thompson5910):

alright

jimthompson5910 (jim_thompson5910):

We need the fractions to have the same and equal denominators.

jimthompson5910 (jim_thompson5910):

So we must get each denominator equal to the LCD

OpenStudy (anonymous):

the lcd of both numbers

jimthompson5910 (jim_thompson5910):

The denominators are 2x and 3x. So the LCD is 6x So multiply the first fraction by 3/3 to get \[\Large \frac{3*3}{3*2x} = \frac{9}{6x}\] and multiply the second fraction by 2/2 to get \[\Large \frac{2*4}{2*3x} = \frac{8}{6x}\]

OpenStudy (anonymous):

thank you i know what to do from here

jimthompson5910 (jim_thompson5910):

So \[\Large \frac{3}{2x}+\frac{4}{3x}\] becomes \[\Large \frac{9}{6x}+\frac{8}{6x}\] and then you can add the numerators and place that over the LCD

jimthompson5910 (jim_thompson5910):

You're welcome. Let me know what you get for your answer.

OpenStudy (anonymous):

17/6x is that correct

jimthompson5910 (jim_thompson5910):

you nailed it, very nice

OpenStudy (anonymous):

thank you for helping

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

its 17/6x :)